OpenStudy (anonymous):

g(f(x))??? f(x)= x-5 g=x squared

OpenStudy (lalaly):

g(f(x)=x-5

OpenStudy (lalaly):

(x-5)^2

OpenStudy (lalaly):

didnt see the squared :$

OpenStudy (anonymous):

\[g \left( f(x) \right)=g \left( x-5 \right)=(x-5)^2\]

OpenStudy (anonymous):

you might want to expand the binomial\[x^2-10x+25\]

OpenStudy (anonymous):

is it x^2(x-5) and do i just destribute?

OpenStudy (anonymous):

the rule of association for the function f, which is x-5 is substituted in the rule from function g, which is x^2

OpenStudy (anonymous):

i think it is x^3-5x^2 correct me if im wrong

OpenStudy (anonymous):

that is incorrect. the answer in two different forms are in my first and second posts:\[(x-5)^2=x^2-10x+25\]

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

hey 1 more thing